How is Gibbs free energy change determined when ΔH and ΔS are available?

ΔG = ΔH - TΔS


Ozone (O3) in the atmosphere can react with nitric oxide (NO):

O3(g) + NO(g) > NO2(g) + O2(g)


Calculate the G for this reaction at 25°C.

(H = -199 kJ/mol, S = -4.1 J/K•mol)


1
Expert's answer
2022-06-03T02:25:03-0400

∆H = -199 kJ

∆S = -4,1 J/K•mol = -0,0041 kJ/K•mol

T = 25°C = 298 K


∆G = ∆H - T∆S

∆G = -199 -( 298 * (-0,0041) ) = -197,7782 kJ


ANSWER: ∆G = -197,7782 kJ ≈ -197,8 kJ